### v2

# Multiple Stochastic Prompt Tuning for Few-shot Adaptation under Extreme Domain Shift

Abstract
Foundation Vision-Language Models (VLMs) like CLIP exhibit strong generalization capabilities due to large-scale pretraining on diverse image-text pairs. However, their performance often degrades when applied to target datasets with significant distribution shifts in both visual appearance and class semantics. Recent few-shot learning approaches adapt CLIP to downstream tasks using limited labeled data via adapter or prompt tuning, but are not specifically designed to handle such extreme domain shifts. Conversely, some works addressing cross-domain few-shot learning consider such domain-shifted scenarios but operate in an episodic setting with only a few classes per episode—limiting their applicability to real-world deployment, where all classes must be handled simultaneously. To address this gap, we propose a novel framework, MIST (Multiple Stochastic Prompt Tuning), for efficiently adapting CLIP to datasets with extreme distribution shifts using only a few labeled examples, in scenarios involving all classes at once. Specifically, we introduce multiple learnable prompts per class to effectively capture diverse modes in visual representations arising from distribution shifts. To further enhance generalization, these prompts are modeled as learnable Gaussian distributions, enabling efficient exploration of the prompt parameter space and reducing overfitting caused by limited supervision. Extensive experiments and comparisons with state-of-the-art methods demonstrate the effectiveness of the proposed framework.
1 Introduction
Foundation Vision-Language Models (VLMs) such as CLIP clip and ALIGN align have significantly advanced computer vision by enabling strong zero-shot generalization across tasks. Trained on large-scale image-text pairs, they learn robust representation spaces. However, such pretraining is costly and impractical to replicate for every new task, prompting growing interest in efficient adaptation using only a few labeled examples.
However, few-shot adaptation of such large-scale models is challenging due to the risk of overfitting and the potential loss of their original pretrained generalization capabilities. To address this, an emerging body of work has explored parameter-efficient fine-tuning techniques, such as prompt tuning coop and adapter tuning clip_adapter . These methods avoid full model fine-tuning by introducing a small set of trainable parameters while keeping the backbone frozen—either by modifying the input space (prompt tuning) or the output layers (adapter tuning). Despite these advances, such approaches primarily focus on standard benchmark datasets with natural images and class semantics (e.g., ImageNet, Caltech101), and largely overlook scenarios involving extreme domain shifts. In real-world applications, downstream datasets can exhibit significant shifts in both visual appearance and label semantics. For example, medical image datasets often feature domain-specific content and class names that do not align with natural image concepts and are typically unavailable for pretraining due to privacy concerns. While recent works have applied CLIP to cross-domain few-shot learning (CDFSL) semantic_guided , they typically adopt a source-free meta-testing setup, adapting to episodes with a few sampled classes (e.g., 5-way) from the target domain. Performance is averaged over many such episodes. However, this approach is computationally expensive and misaligned with real-world settings, where all target classes are present simultaneously.
In this work, we propose a novel prompt learning framework, MIST (Multiple Stochastic Prompt Tuning), for adapting CLIP to a more realistic setting, where target datasets exhibit significant domain and semantic shifts, and only a few labeled examples from all classes are available simultaneously. We first observe that extreme distribution shifts can lead to fragmented visual representations, forming separate and inconsistent clusters in the embedding space (Fig. 2). Moreover, large number of classes with semantic shifts (different class labels) can cause multiple class features to cluster together, resulting in ambiguous decision boundaries. To address these challenges, we introduce multiple learnable prompts per class, enabling better modeling of multi-modal feature distributions. Further, instead of directly optimizing prompt weights, we represent each prompt as a Gaussian distribution with learnable mean and variance, promoting diverse and well-separated representations while mitigating overfitting through efficient exploration of the prompt space. The key contributions of this work are summarized below:
  1. 1.
    We propose a novel framework for few-shot adaptation of CLIP to realistic scenarios involving extreme domain and label semantic shifts, with all target classes present simultaneously.
  2. 2.
    We analyze limitations of existing prompt tuning methods for few-shot setting, under severe distribution shifts.
  3. 3.
    We propose MIST, a novel prompt tuning framework that uses multiple class-specific prompts to model multimodal visual feature distributions.
  4. 4.
    We further represent each prompt as a learnable Gaussian distribution, enabling better generalization and reducing overfitting in low-data regimes.
  5. 5.
    Extensive experiments demonstrate that MIST outperforms state-of-the-art methods across multiple challenging benchmarks.
2 Related Work
Here, we briefly discuss the related work in literature.
Vision-language foundation models. Recently, foundation Vision-Language Models (VLMs) clip ; align ; blip have shown strong zero-shot generalization by learning aligned visual-textual representations from web-scale image-text pairs. However, their performance can degrade on domain-specific tasks with distribution shifts or rare class semantics not seen during pretraining. As retraining such large models for new tasks is impractical, recent efforts focus on efficient adaptation using limited labeled samples from the target domain.
Few-shot adaptation of VLMs. Adapting these large-scale models to downstream tasks with few labeled training data is often challenging, due to the risk of overfitting. Efficient transfer learning methods like prompt tuning coop ; maple ; promptsrc or adapter tuning clip_adapter ; tip_adapter address this issue by optimizing only a few parameters added to these models, either in the input space or in the intermediate or output layers. For instance, CLIP-Adapter clip_adapter trains a classifier on the visual output features to modify the visual feature space, and Tip-Adapter tip_adapter stores the few-shot image prototypes, which are used to compute similarity with the test samples to guide the visual encoder. In prompt-tuning, CoOp coop trains a few prompt vectors appended to the classname text, keeping the CLIP encoders frozen. MaPLe maple proposes training prompts in both the textual and visual branches to improve multimodal alignment, while PromptSRC promptsrc further enhances performance by distilling knowledge from the frozen CLIP model. tcp incorporates class description features into the text encoder during prompt tuning to improve discriminability of the classifier. proda tunes text prompts using a pool of diverse prompts for each class, while bayesian_pl adds learnable gaussian noise to each token of the learnable text prompt, to improve generalization. These works mainly focus on standard image datasets with natural images and class semantics, often overlooking more realistic deployment scenarios where the target datasets may exhibit substantial domain shifts or unfamiliar, specialized label semantics. Some recent works have employed CLIP for the cross-domain few-shot learning (CDFSL) setup, which incorporates these challenges semantic_guided ; prompt_free . However, these methods take the episodic paradigm, which samples fixed number of classes in each episode, which is often unrealistic in real-world deployment.
Stochastic neural networks. Standard neural networks train weights deterministically as point-estimates. Contrarily, Bayesian Neural Networks bnns ; weight_uncertainty model the weights as probability distributions, making them useful in handling uncertainty in predictions as well as learning robust representations. Stochastic classifiers have been explored in UDA star , person re-identification re-id , incremental learning s3c and DG stylematch in prior literature. To the best of our knowledge, this is the first work which explores stochastic classifiers for few-shot adaptation of VLMs under extreme domain shifts.
3 Problem Formulation
Given few labeled training examples from a target dataset, the task is to adapt the CLIP model efficiently to this data. Formally, we have a support set $\mathcal{S}=\{(X_{i},y_{i})\}^{C\times k}_{i=1}$ from the target dataset containing $k$ samples from all the $C$ classes simultaneously. Here, $y_{i}\in\{0,1\}^{C}$ is the corresponding ground truth label, and $k=\{1,2,4,8,16\}$, denotes the number of shots. The evaluation is done on the full test set. Here, in addition to the few-shot problem, the target dataset contains significant domain and label semantic shift from natural image datasets.
3.1 Preliminaries
Here, we briefly describe the process of classifying images with CLIP and the base network used in this work for completion. Let us denote the CLIP text and image encoders as $\mathcal{F}_{t}$ and $\mathcal{F}_{v}$ respectively. The input image $Xv=\mathbb{R}^{C\times H\times W}$ is broken up into patches $\{e_{CLS},e_{1},e_{2},...,e_{M}\}$ and fed into the image encoder to extract the image embedding $z_{v}=\mathcal{F}_{v}(X_{v})$. Similarly, the text input (typically of the form “A photo of a [CLS]”) is tokenized in the form $X_{t}=\{t_{SOS},t_{1},t_{2},...,t_{CLS},t_{EOS}\}$ and fed into the text encoder to get the text embedding $z_{t}=\mathcal{F}_{t}(X_{t})$. During zero shot classification, the class text embeddings are matched with the image as follows: $\frac{exp(<z_{v},z_{t}>/\tau)}{\sum\limits_{i=1}^{C}exp(<z_{v},z_{t_{i}}>/\tau)}$, where $C$ is the number of classes and $\tau$ is the temperature constant. The class with highest similarity is the prediction.
Base Network of MIST: In MIST, we employ a multimodal prompt learning strategy, where learnable prompt vectors are appended to the image and textual input branches. Specifically, let the learnable text prompt vectors be denoted as $\theta_{t}=\{\theta_{t_{1}},\theta_{t_{2}},...,\theta_{t_{m}}\}$ and the visual prompts as $\theta_{v}=\{\theta_{v_{1}},\theta_{v_{2}},...,\theta_{v_{m}}\}$. These are appended to the input text and image patches to form the modified inputs as: $\tilde{X}_{t}=\{t_{SOS},\theta_{t_{1}},...,\theta_{t_{m}},t_{1},t_{2},...,t_{CLS},t_{EOS}\}$ and $\tilde{X}_{v}=\{e_{CLS},\theta_{v_{1}},...,\theta_{v_{m}},e_{1},e_{2},...,e_{M}\}$ respectively. The extracted feature embeddings from the CLIP encoders are now $\tilde{z}_{t}=\mathcal{F}_{t}(\tilde{X}_{t})$ and $\tilde{z}_{v}=\mathcal{F}_{v}(\tilde{X}_{v})$. Here, the trainable textual prompts are passed through a learnable projection layer $f_{\phi}$ to obtain the visual prompts, i.e., $\theta_{v}=f_{\phi}(\theta_{t})$. Along with adding prompts to the inputs, we also adopt a deep prompting approach maple ; promptsrc , where learnable prompt vectors are attached after every transformer block. When adapting to a downstream task, these multimodal prompts are trained in an end-to-end manner, keeping the CLIP encoders frozen.
4 The Proposed Framework
The goal is to efficiently adapt the large-scale pre-trained CLIP using very few samples from all the classes simultaneously, under extreme domain and semantic shifts. To this end, we present MIST (Fig. 3), which augments the base network with two novel modules to tackle these challenges: (i) Stochastic prompt learning to mitigate the risk of overfitting due to very few training examples and (ii) Multiple prompts to address the implicit unimodal assumption regarding the class data distributions, which we descrie below.
Figure 1: Effect of deterministic vs stochastic prompt learning on the EuroSAT dataset with 1-shot per class. The projected class-specific image features are spread in fixed regions for stochastic learning, implicitly learning well-separated decision boundaries.
Method EuroSAT ISIC
1-Shot
Deterministic Prompt Tuning 73.30 27.50
Stochastic Prompt Tuning ($\mu,\sigma^{*}$) 73.43 30.67
Stochastic Prompt Tuning ($\mu^{*},\sigma^{*}$) 67.40 22.67
8-Shots
Deterministic Prompt Tuning 86.80 46.53
Stochastic Prompt Tuning ($\mu,\sigma^{*}$) 86.73 50.80
Stochastic Prompt Tuning ($\mu^{*},\sigma^{*}$) 88.03 50.93
Table 1: Stochastic prompt learning with two different sampling techniques on EuroSAT and ISIC datasets. The fixed mean approach ($\mu,\sigma^{*}$) performs better in low shots, while sampling from a fully learnable distribution ($\mu^{*},\sigma^{*}$) performs better in higher shots.
Figure 2: t-SNE visualization of image embeddings from MaPLe maple (left) and TCP tcp (right) for the EuroSAT dataset. The classwise data distribution violates the unimodal assumption, due to high interclass similarity and intraclass variations.
Figure 3: Our proposed MIST framework. We append two sets of prompts to the classnames, one sampled from a fixed mean ($\bar{\mu}_{0},\sigma_{0}$) and the other from a fully learnable Gaussian distribution ($\mu_{1},\sigma_{1}$). $f_{\phi}$ projects the text prompts to visual prompts. The loss term $\mathcal{L}_{mp}$ trains the distribution parameters ($\mu_{1},\sigma_{0},\sigma_{1}$) and $f_{\phi}$ such that the image embedding is assigned to the closest text prototype of its respective class. The $\mathcal{L}_{reg}$ term prevents the two prompts from collapsing by enforcing diversity in the class-wise prompt training.
4.1 Stochastic Prompt Learning
Although existing prompt tuning methods optimize a few learnable parameters appended to the inputs, the limited availability of training samples makes large-scale models like CLIP prone to overfitting as observed by promptsrc . Additionally, when all target classes are present during adaptation, domain-specific or unfamiliar classnames can cause the class-specific features to lie closer together in the representation space, making inter-class boundaries harder to distinguish. To address these issues, we explore a novel strategy using stochastic prompt learning, where we model the prompts with learnable distributions.
The main idea is to learn a distribution over the prompt parameters instead of optimizing them as point-estimates as is the standard practice. This helps to mitigate the uncertainty arising from scarce data, since each distinct sampled weight from this learnable distribution forms diverse decision boundaries for the few shot data, by allowing a richer exploration of the prompt parameter space. This provides an implicit regularization to the model without additional loss functions leading to more robust decision boundaries for the few-shot training samples. Specifically, we sample the text prompt weights from a Gaussian distribution $\mathcal{N}(\mu,\sigma)$, parameterized by mean $\mu$ and variance $\sigma$ as follows: $\theta_{t}\sim\mathcal{N}(\mu,\sigma)$. After every iteration, we backpropagate the loss to the learnable $\mu$ and $\sigma$ parameters. For end-to-end training, we use the Gaussian reparameterization trick var_auto_encoder as follows:
$\theta_{t}=\mu+\mathcal{N}(0,I)\odot\sigma$ (1)
We observe from Fig. 1 that in the deterministic case, passing the same examples through the trained model always projects them to the same points in the feature space. In contrast, the projections in the stochastic scenario spreads over a broader region, due to sampling of weights from the learned distribution. This variability implicitly encourages a margin between the class-specific features, resulting in more discriminative decision boundaries.
To verify this, we consider two distinct strategies for sampling the text prompt parameters. First, we fix the mean of the Gaussian to the standard prompt “A photo”, keeping the variance learnable. In the second case, both the mean and variance are learnable. The results on two datasets bscdfsl are shown in Table 1. We observe that in low-shot setting, the first approach outperforms the second, while an opposite trend is seen in the higher-shot setting. This suggests that with very few training data (1-shot), directly optimizing the parameters of a distribution is challenging, but exploring variations around a standard prompt improves performance. On the contrary, learning the full distribution with more data outperforms the fixed mean approach. Hence, the two strategies complement each other in facilitating an efficient coverage of the prompt parameter space.
4.2 MIST: Multiple Stochastic Prompt Tuning
Since CLIP is pretrained on web-scale data encompassing mainly natural images clip , the image encoder struggles to learn robust classwise features when faced with extreme domain and semantic shifts in the target dataset. In addition, intraclass diversity and interclass similarity in the images due to presence of many classes, results in disjoint clusters in visual features from the same class. Consider an illustrative example of the EuroSAT dataset containing satellite images of various terrains. Here, distinct classes like “Highway” and “River” can have similar visual representatives in the few-shot training data, at the same time featuring diverse visuals from the same classes as shown in Fig. 2. Existing prompt tuning approaches represent each class with single learnable prompts, implicitly assuming that classwise visual features form single clusters. However, such a strategy is insufficient to represent the disjoint visual clusters, which may result from such challenging settings. To illustrate this, we consider two representative prompt-tuning methods, MaPLe maple and TCP tcp and show their t-SNE visualizations in Fig. 2. We observe that the unimodal assumption is violated and the image embeddings of each class form multiple modes in the representation space.
To address this, inspired from imp ; mixtfsl , we introduce a multiple prompt learning approach, where we represent each class with multiple prompt vectors. However, incorporating too many learnable prompts for each class is not desirable since: (i) Each class contains few training samples, and many text embeddings per class could lead to overfitting on individual datapoints, resulting in loss of class level representations, and (ii) it introduces additional learnable parameters and thus more computational overhead. As a balanced approach, we represent each class with two prompt vectors. Formally, let the embeddings corresponding to the two prompts for a particular class $CLS_{k}$ be denoted as $\tilde{z}_{t}^{i}$, where, $i$ = 1, 2. Here, $\tilde{z}_{t}^{i}=\mathcal{F}_{t}(\tilde{X}_{t}^{i})$, where, $\tilde{X}_{t}^{i}=\{t_{SOS},\theta_{t}^{i},t_{1},...,t_{CLS_{k}},t_{EOS}\}$ is the $i^{th}$ text prompt for the class $CLS_{k}$. $\theta_{t}^{i}=\{\theta_{t1}^{i},\theta_{t2}^{i},...,\theta_{tm}^{i}\}$ denotes the $i^{th}$ set of learnable prompt vectors for that particular class.
To simultaneously represent the underlying multimodal class distribution and mitigate overfitting, we stochastically model the parameters of the two prompts for each class as described in Sec 4.1. Specifically, we incorporate the fixed mean, learnable variance approach on the parameters of the first prompt, simultaneously learning a full Gaussian distribution over the parameters of the second prompt:
$\theta_{t}^{1}\sim\mathcal{N}(\bar{\mu}_{0},\sigma_{0})$ (2)
$\theta_{t}^{2}\sim\mathcal{N}(\mu_{1},\sigma_{1})$
Here, $\bar{\mu}_{0}$ is a fixed vector corresponding to the text embedding of “A photo”, and $\mu_{1}$, $\sigma_{0}$, $\sigma_{1}$ are learnable parameters. Now, we describe the training process of MIST.
MIST Training: For a particular image embedding $\tilde{z}_{v}$, we first find the closest text embedding of its respective class after every iteration, based on cosine similarity as follows:
$i^{*}=\operatorname*{argmax}_{i\in\{1,2\}}\,\,\,sim(\tilde{z}_{t}^{i},\tilde{z}_{v})$ (3)
where, $sim(a,b)=\frac{a\cdot b}{\lVert a\rVert\lVert b\rVert}$ denotes the cosine similarity. During training, the image embedding is assigned to its closest prompt embedding by minimizing the following loss function:
$\mathcal{L}_{mp}=-log\,\Big{(}\frac{exp(sim(\tilde{z}_{t}^{i^{*}},\tilde{z}_{v}))}{\sum\limits_{j=1}^{2C}exp(sim(\tilde{z}_{t}^{j},\tilde{z}_{v}))}\Big{)}$ (4)
where, $C$ is the number of classes, and $sim(\cdot)$ denotes the cosine similarity. To prevent the image embedding $\tilde{z}_{v}$ from being always assigned to a single text prompt, and encourage diversity when training the two prompts within the same class, we minimize an additional regularization term to increase the cosine similarity of the image embedding to the centroid of the two text embeddings of its corresponding class:
$\mathcal{L}_{reg}=-sim(\tilde{z}_{v},\tilde{z}_{t}^{c})$ (5)
where, $\tilde{z}_{t}^{c}=\frac{1}{2}(\tilde{z}_{t}^{1}+\tilde{z}_{t}^{2})$ is the centroid of the prompt embeddings for the corresponding class and $sim(\cdot)$ represents the cosine similarity. Thus, the final objective function is:
$\mathcal{L}_{total}=\mathcal{L}_{mp}+\mathcal{L}_{reg}$ (6)
This loss function is used to optimize the Gaussian parameters $\mu_{1}$, $\sigma_{0}$ and $\sigma_{1}$, as well as the projection layers as:
$\mu_{1}^{*},\sigma_{0}^{*},\sigma_{1}^{*},\phi^{*}=\operatorname*{argmin}_{\mu_{1},\sigma_{0},\sigma_{1},\phi}\mathbb{E}_{(X,y)\sim D_{tgt}}\mathcal{L}_{total}(X,y)$ (7)
Method EuroSAT ISIC PDisease ChestX Average
1-shot
CoOp (IJCV’22) $51.87$ $22.77$ $24.73$ 22.83 $30.55$
TaskRes (CVPR’23) $64.67$ $19.70$ $36.57$ $10.97$ $32.98$
MaPLe (CVPR’23) $73.30$ $27.50$ $51.53$ $14.60$ $41.73$
PromptSRC (ICCV’23) $73.23$ $21.97$ 55.03 $14.37$ $41.15$
CLAP (CVPR’24) $61.46$ $26.61$ $47.22$ $15.94$ $37.81$
TCP (CVPR’24) $64.30$ $27.80$ $49.37$ $14.93$ $39.10$
MIST (Ours) 77.90 34.40 $50.27$ $17.10$ 44.92
2-shot
CoOp (IJCV’22) $66.00$ $21.87$ $37.97$ $14.43$ $35.07$
TaskRes (CVPR’23) $68.83$ $23.13$ $39.27$ $10.83$ $35.52$
MaPLe (CVPR’23) $78.07$ $31.90$ $67.17$ $16.27$ $48.35$
PromptSRC (ICCV’23) $79.53$ $29.47$ $68.07$ $12.70$ $47.44$
CLAP (CVPR’24) $70.63$ $34.79$ $60.13$ 16.43 $45.50$
TCP (CVPR’24) $70.37$ 36.87 $62.63$ $15.63$ $46.38$
MIST (Ours) 81.57 $36.37$ 69.60 $13.90$ 50.36
4-shot
CoOp (IJCV’22) $66.53$ $25.00$ $42.67$ $17.93$ $38.03$
TaskRes (CVPR’23) $72.40$ $21.40$ $39.35$ $10.27$ $35.86$
MaPLe (CVPR’23) $84.03$ $37.17$ $77.07$ 19.73 $54.50$
PromptSRC (ICCV’23) $85.23$ $37.63$ $78.70$ $15.17$ $54.18$
CLAP (CVPR’24) $76.43$ $34.37$ $65.11$ $18.98$ $48.72$
TCP (CVPR’24) $76.77$ $37.37$ $67.97$ $17.07$ $49.80$
MIST (Ours) 85.93 40.90 79.67 $18.67$ 56.29
8-shot
CoOp (IJCV’22) $76.53$ $38.27$ $60.50$ $14.60$ $47.48$
TaskRes (CVPR’23) $74.63$ $34.30$ $57.77$ $12.57$ $44.82$
MaPLe (CVPR’23) $86.80$ $46.53$ $84.47$ $14.17$ $57.99$
PromptSRC (ICCV’23) $88.37$ $42.47$ $86.80$ $14.93$ $58.14$
CLAP (CVPR’24) $76.85$ $42.81$ $74.62$ $14.97$ $52.31$
TCP (CVPR’24) $79.03$ $46.57$ $76.33$ $14.97$ $54.23$
MIST (Ours) 88.63 52.70 87.47 16.50 61.33
16-shot
CoOp (IJCV’22) $82.83$ $43.40$ $69.90$ 18.80 $53.73$
TaskRes (CVPR’23) $79.90$ $38.10$ $69.40$ $12.87$ $50.07$
MaPLe (CVPR’23) $92.80$ $55.53$ $89.93$ $13.90$ $63.04$
PromptSRC (ICCV’23) $92.55$ $55.17$ $91.40$ $14.83$ $63.49$
CLAP (CVPR’24) $82.96$ $49.43$ $78.27$ $17.47$ $57.03$
TCP (CVPR’24) $84.93$ $52.83$ $80.63$ $16.53$ $58.73$
MIST (Ours) 93.57 60.30 91.73 $14.77$ 65.09
Table 2: Performance comparison (average accuracy (%) over 3 seeds) of the proposed MIST with the state-of-the-art approaches for $k=1,2,4,8,16$ shots from each class.
Inference: After learning the parameters of the distribution, during inference, we can sample weights for the two text prompts as follows: $\theta_{t}^{1}\sim\mathcal{N}(\bar{\mu}_{0},\sigma_{0}^{*})$ and $\theta_{t}^{2}\sim\mathcal{N}(\mu_{1}^{*},\sigma_{1}^{*})$. For each class, we take the maximum logit among the two text prompts as the output prediction for that class.
5 Experimental Results
Here, we extensively evaluate the proposed framework and compare it with the state-of-the-art approaches.
Datasets used: For the target datasets, we consider the BSCDFSL bscdfsl benchmark, which is collected from real-world settings, and consists of four datasets, namely EuroSAT eurosat , ISIC isic , Plant Disease plant_disease and ChestX chestx . These datasets cover a varying spectrum of domain shifts, along with specialized classnames, encompassing satellite, agricultural and medical images. For training, we consider few samples ($k=1,2,4,8,16$) randomly selected from all the classes together and then evaluate the trained model on the full test set of all the datasets. The final accuracy is taken as the average over 3 different seeds.
Implementation details: We employ CLIP ViT-B/16 as the backbone similar to MaPLe maple . The learnable context length of the text and vision inputs are set as $2$, and deep prompts are incorporated upto a depth of $9$. The model is trained using SGD optimizer for $150$ epochs with a learning rate of $0.0035$ and a batch size of $4$. All experiments are conducted on a NVIDIA RTX A5000 GPU.
5.1 Comparison with state-of-the-art methods
To validate the effectiveness of our approach, we compare our proposed MIST with several recent CLIP-based efficient transfer learning methods for varying number of shots. Specifically, we compare with 1) CoOp coop and TCP tcp , which employ prompt tuning on the text branch; 2) MaPLe maple and PromptSRC promptsrc which utilize a multimodal prompt tuning approach; 3) TaskRes taskres where task-specific adapters are tuned keeping the base text classifier fixed; 4) CLAP clap uses a linear-probing approach and mainly addresses the absence of validation sets in FSL.
For fair comparison, we run all the methods (using the official, publicly available codes) on the ViT-B/16 backbone and report the results in Table 2. We list some observations below:
(i) Among the competing methods, MaPLe and PromptSRC achieves the highest performance on average, closely followed by TCP. Their improved performance suggests the effectiveness of multimodal prompt tuning in handling distribution shifts over text prompt tuning, which was also observed in maple . This observation is further supported by our results;
(ii) Although, CLAP is a recent approach, it mainly focuses on the validation problem of FSL. The reduced performance of CLAP highlights the limitations of linear probing, which does not utilize the text information for handling significant semantic and domain shifts;
(iii) As the number of shots increases, the multimodal prompt tuning approaches like MaPLe, PromptSRC and MIST outperforms other methods by larger margins, suggesting that training more parameters is more effective for higher shots.
(iv) Although all the methods show similar performance on the ChestX dataset, their overall accuracies are extremely low due to the large domain shift. However, text prompt tuning methods perform slightly better than the multimodal counterparts. This maybe because ChestX contains greyscale images, where additional prompt tuning in the vision branch degrades the performance. Overall, our proposed MIST framework outperforms the other methods significantly, giving consistent average gains of $3.19\%$, $2.01\%$, $1.79\%$, $3.19\%$, $1.60\%$ on $k=1,2,4,8,16$ shots respectively, over the best performing methods. The significant improvement for the 1-shot case highlights the effectiveness of our approach in mitigating overfitting in extremely low data scenarios.
Dataset MaPLe PromptSRC TCP MIST (Ours)
EuroSAT $72.90$ $74.10$ $57.80$ 76.90
ISIC $14.80$ $16.10$ $13.13$ 16.73
Table 3: Performance comparison (%) of MIST with state-of-the-art methods on the class-imbalanced setting, with varying data samples from each class.
Figure 4: Generalization to classes: The class-wise accuracies are sorted and divided into 2 bins. Our proposed MIST outperforms the other methods in both the bins and also increases the worst-class accuracy for the same seed in the challenging 1-shot setting.
5.2 Additional Analysis
Here we perform additional analysis and ablation studies to further validate our proposed framework. For the analysis, we compare with MaPLe, PromptSRC since they use multimodal prompts and are better suited for this task and TCP, since it is the state-of-the-art prompt tuning approach on CLIP.
1) Class-imbalanced learning: Here, we explore an even more challenging scenario, where the number of labeled examples may vary across classes, reflecting real-world datasets. The standard few-shot settings in literature assume an idealistic scenario where each class has exactly $k$ training samples, overlooking the effect of class imbalance. To create such a setting, we perform data sampling in a cyclic manner, e.g., we take $\{1,2,4,8,1,2,$...\} from each class of the target dataset for training. The model is then evaluated on the entire test set. The results on two representative datasets, EuroSAT and ISIC in Table 3 shows that the proposed MIST outperforms the other methods even under class-imbalanced conditions, highlighting its effectiveness.
2) Sensitivity to training samples: The representation learning capability of a model is largely influenced by the few sampled training examples across classes. However, a robust model should ideally achieve a lower variance across different sampling strategies. To account for this, all the reported results are accuracies averaged over 3 different random seeds. Here, we further report the variance across the 3 seeds for the EuroSAT and ISIC datasets in Table 4. We observe that the proposed MIST not only achieves a higher accuracy, but also shows a much lower variance, highlighting its robustness to different sampling techniques.
3) Generalization to all classes: In practical scenarios, the overall accuracy is often not a reliable metric to understand the model’s ability to represent difficult classes. Here, we study the effectiveness of our multiple stochastic prompt-tuning approach in learning generalized class boundaries and modeling all the complex class distributions.
EuroSAT ISIC
1-shot MaPLe $73.30$ $\pm 3.84$ $27.50$ $\pm 10.19$
PromptSRC $73.23$ $\pm 3.75$ $21.97$ $\pm 6.09$
TCP $64.30$ $\pm 3.24$ $27.80$ $\pm 6.55$
MIST (Ours) $77.90$ $\pm\textbf{2.63}$ 34.40 $\pm\textbf{3.56}$
2-shots MaPLe $78.07$ $\pm 5.87$ $31.90$ $\pm 6.08$
PromptSRC $79.53$ $\pm 2.76$ $29.47$ $\pm 6.50$
TCP $70.37$ $\pm 2.31$ $36.87$ $\pm 8.29$
MIST (Ours) $81.57$ $\pm\textbf{1.84}$ $36.37$ $\pm\textbf{3.96}$
Table 4: MIST exhibits significantly lower variance across three different random seeds compared to other approaches.
Specifically, we first sort the class-wise accuracies in ascending order. The classes are then divided into two bins in this order to highlight the gain in accuracy in both the lower as well as the higher bin. The comparison with the other methods are shown in Figure 4 for one random seed (same for all methods). We observe that the proposed MIST improves the accuracies in both the bins, while also increasing the worst-class accuracy, which indicates that MIST learns more generalized class representations.
4) Ablation Study: Our proposed MIST framework models the two text classifiers using two distinct Gaussian sampling techniques as described earlier. Here we analyze the effectiveness of each of the proposed components in Table 5. Our base method is the multimodal prompt tuning framework, MaPLe. Introducing stochasticity to this single prompt (fixed mean) tuning approach by sampling from a learnable Gaussian distribution mitigates overfitting as described earlier, and improves performance. Introducing the second learnable prompt sampled from a fully learnable distribution without the $\mathcal{L}_{reg}$ term increases performance in EuroSAT, but reduces for ISIC. This can happen due to collapse of the two classifiers without the regularization term mixtfsl ; deepfake . Finally, adding the regularization term $\mathcal{L}_{reg}$ outperforms the baseline significantly for both the datasets.
Method EuroSAT ISIC
Base Network $73.30$ $27.50$
+ Stochastic Prompt Learning ($\mu$, $\sigma^{*}$) $73.43$ $30.67$
+ Multiple Stochastic Prompting $74.27$ $27.00$
+ Multiple Stochastic Prompting + $\mathcal{L}_{reg}$ (MIST) 77.90 34.40
Table 5: Ablation study (1-shot): All the proposed components collectively enhance the overall performance.
Figure 5: Effect of prompt length (left) and number of text prompts per class (right) for EuroSAT data (1-shot).
5) Number of text prompts & Prompt Length: MIST utilizes two prompts per class sampled from learnable Gaussian distributions. Here we study the effect of adding more text prompts. For this, we keep one prompt with a fixed mean, and the others sampled from fully learnable distributions. From Figure 5 (right), we observe that the performance starts decreasing after three prompts, suggesting overfitting from the increasing number of learnable parameters. Further, addition of more classifiers introduces increased computational overhead and longer training time. The effect of increasing the number of learnable prompt vectors is illustrated in Figure 5 (left). Here also, the accuracy decreases after a point, indicating overfitting in the few-shot setting. We used 2 learnable prompts for all the experiments.
Qualitative Results: We illustrate the inherent challenges of these datasets in Figure 6, along with some of the predictions from the proposed MIST framework.
Figure 6: Qualitative results: From left to right, shows predictions on EuroSAT, ISIC, Plant Disease and ChestX respectively. Green denotes correct while red denotes incorrect predictions.
Limitations. While MIST outperforms state-of-the-art methods across all datasets, its performance slightly drops on the grayscale ChestX dataset, likely because of the additional visual prompts. In such cases, methods relying solely on textual prompts may prove more effective.
6 Conclusion
In this work, we propose a novel framework, MIST for adapting foundation VLMs like CLIP to realistic few-shot scenarios characterized by extreme domain and label semantic shifts. Motivated by the limitations of existing parameter efficient fine-tuning approaches, we incorporate multiple text prompts per class, modeled by distinct learnable Gaussian distributions to represent the inherent multimodal class distributions as well as mitigate overfitting. Extensive experiments on multiple benchmarks as well as additional analysis show the effectiveness of our proposed approach compared to state-of-the-art methods.